OpenStudy (anonymous):

exponents (-t)(-t)^2(-t^4) please tell me how you get the answer, not just the answer please what is the rule for exponents outside of the parentheses, like (-t)^2 and there's no exponent inside the parentheses?

OpenStudy (anonymous):

You have to take the exponent for each part inside. So (t^2) is just t^2. (2t)^2 would be 2^2 times t^2 or 4t^2. For yours, (-t)*(t^2)(-t^4). The second t is like multiplying -1*-1=1 or -2*-2=4. 4 is really just 2^2 so with variables we can leave it as t^2. (-t)*(t^2) is -t^3 and then multiplied by (-t^4) you'd have (-t^3)*(-t^4)=t^7

OpenStudy (anonymous):

thank you! so, we multiply the base number, but we add the exponents, is that correct?

OpenStudy (anonymous):

correct. so what would x^2 * x^2 equal?

OpenStudy (anonymous):

x^4?

OpenStudy (anonymous):

exactly. You can see that it would also be (x^2)^2 if you square x^2, you get x^4. Good!

OpenStudy (anonymous):

sorry, but can you clarify one more thing for me? so the second variable for this equation, (-t)(-t)^2(-t^4) (-t)^2= t^2, it becomes positive if it were, (-t)3, it remains as -t^3, it wouldn't become positive right?